OpenStudy (anonymous):

Evaluate the fallowing integral where in case C is the circle \[\left| z \right|\] =3, counterclockwise \[\int\limits_{C}^{}zdz/(z ^{2}+1)\]

OpenStudy (anonymous):

@experimentX do u know cauchy teorem bro ?

OpenStudy (experimentx):

no ... man ... i'll see

OpenStudy (anonymous):

there are a lot of cauchy theorems which one .?

OpenStudy (anonymous):

about contour i guess

OpenStudy (experimentx):

can you post link ... or on wikipedia?? i'll see in a while

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Methods_of_contour_integration @experimentX i think thats the one

OpenStudy (anonymous):

Specifically the residue theorem: http://en.wikipedia.org/wiki/Residue_theorem

OpenStudy (anonymous):

isn't Cauchy-Goursat theorem ?!

OpenStudy (anonymous):

residue herp_derp is right :/ i am wrong

OpenStudy (anonymous):

The circle encloses two simple poles (one at i, the other at -i), so their residues can be calculated by:\[R=\lim_{z \rightarrow c} (z-c) f(z)\]For c=i and c=-i. The contour integral is simply 2πi times the sum of these residues, by Cauchy's residue theorem.

OpenStudy (anonymous):

Both residues are 1/2, so the integral is simply 2πi.

OpenStudy (experimentx):

looks like i should go back to study and catch up with you guys!!

OpenStudy (anonymous):

@Herp_Derp i could not understand :/

OpenStudy (anonymous):

my english is not very well :/ maybe you can show some calculation ha ?

OpenStudy (anonymous):

? @Herp_Derp

OpenStudy (anonymous):

\[\int_C \frac{z~\mathrm{d}z}{(z+i)(z-i)}=2\pi i\sum_k \mathrm{Res}(f,a_k)\]\[f(z)=\frac{z}{(z+i)(z-i)},~\{a_k\}=\{-i,i\}=\{c:f~\mathrm{has~a~singularity~at~c}\}\]\[\mathrm{If~}f(z)=\frac{g(z)}{z-c}~\mathrm{for~some~g~which~is~analytic~around~}c,~\mathrm{then~Res}(f,c)=\lim_{z \rightarrow c}(z-c)f(z)=g(c).\]\[\mathrm{Therefore~Res}(f,-i)=\mathrm{Res}(f,i)=1/2~\mathrm{and}~\int_C \frac{z~\mathrm{d}z}{z^2+1}=2\pi i\]